My basic thought is this;
x - x(y) = z
Where x represents my gross pay, y represents an unknown percentage and z represents my net pay.
For example: I make $15.00 per hour and work 40 hours per week, so my gross pay is $600.00. However, my net pay is $531.00. How do I find what percentage comes out of my gross pay to give me my net pay? (Hint: y = 11.5%)
In the past I have found it by trial and error. What I'm looking for is the formula that determines this. Can you help me?
Thank you,
David Iannarelli
thegiantslayer@gmail.com

Answer by nerdybill(7384)

all you have to do is solve for y:
x - x(y) = z
x(1-y) = z
(1-y) = z/x
-y = z/x - 1
y = 1 - z/x
.
that is, percent reduction
1 - "net pay"/"gross pay"
.
1 - 531/600
1 - .885
.115
.
Multiply the above to give the "percentage":
.115 * 100 = 11.5%